# Grandmaster-Level Chess Without Search

### v1

@@ Abstract
The recent breakthrough successes in machine learning are mainly attributed to scale: namely large-scale attention-based architectures and datasets of unprecedented scale. This paper investigates the impact of training at scale for chess. Unlike traditional chess engines that rely on complex heuristics, explicit search, or a combination of both, we train a 270M parameter transformer model with supervised learning on a dataset of 10 million chess games. We annotate each board in the dataset with action-values provided by the powerful Stockfish 16 engine, leading to roughly 15 billion data points. Our largest model reaches a Lichess blitz Elo of 2895 against humans, and successfully solves a series of challenging chess puzzles, without any domain-specific tweaks or explicit search algorithms. We also show that our model outperforms AlphaZero’s policy and value networks (without MCTS) and GPT-3.5-turbo-instruct. A systematic investigation of model and dataset size shows that strong chess performance only arises at sufficient scale. To validate our results, we perform an extensive series of ablations of design choices and hyperparameters.
@@ 1 Introduction
One of the most iconic successes of AI is IBM’s Deep Blue (Campbell et al., 2002) defeating the world chess champion Garry Kasparov in 1997. This was widely seen as the first major demonstration that machines are capable of out-competing humans in intellectual domains that require sophisticated rational reasoning and strategic planning—feats of intelligence that were long believed to be exclusive to humans. Deep Blue was an expert system that combined an extensive database of chess knowledge and heuristics with a strong tree search algorithm (alpha-beta pruning). Almost all modern and much stronger chess engines follow a similar recipe, with Stockfish 16 currently being the world’s strongest (publicly available) engine. Notable exceptions are DeepMind’s AlphaZero (Silver et al., 2017), which uses search and self-taught heuristics but no human chess knowledge, and its open-source replication Leela Chess Zero, which currently often comes in as a close second in chess computer competitions (Haworth and Hernandez, 2021).
Recent breakthroughs in scaling up AI systems have resulted in dramatic progress in cognitive domains that remained challenging for earlier-generation systems like Deep Blue. This progress has been driven by general-purpose techniques, in particular (self-) supervised training on expert data with attention-based architectures (Vaswani et al., 2017) applied at scale, resulting in the development of LLMs with impressive and unexpected cognitive abilities like OpenAI’s GPT series (Brown et al., 2020; OpenAI, 2023), the LLaMA family of models (Touvron et al., 2023a, b), or Google DeepMind’s Chinchilla (Hoffmann et al., 2022) and Gemini (Anil et al., 2023). However, it is unclear whether the same technique would work in a domain like chess, where successful policies typically rely on sophisticated algorithmic reasoning (search, dynamic programming) and complex heuristics. Thus, the main question of this paper is: Is it possible to use supervised learning to obtain a chess policy that generalizes well and thus leads to strong play without explicit search?
To study this question we apply the success recipe of general supervised training at scale to chess (see Figure 1). We use a standard attention-based architecture and a standard supervised training protocol to learn to predict action-values (corresponding to win-percentages) for chess boards. The strength of the resulting chess policy thus depends entirely on the strength of the underlying action-value predictor. To get a large corpus of “ground-truth” action-values we use Stockfish 16 as an oracle to annotate millions of board states obtained from randomly drawn games on lichess.org, which are mostly played by humans varying significantly in playing strength. As we will show this leads to a strong, grandmaster-level chess policy (Lichess blitz Elo 2895 against humans), driven by a modern transformer to predict action-values without any explicit search. This policy outperforms GPT-3.5-turbo-instruct (and, therefore, GPT-4 (Carlini, 2023)) and AlphaZero’s policy and value networks, which reach Elo ratings of 1755, 1620, and 1853, respectively. Therefore, our work shows that it is possible to distill a good approximation of Stockfish 16 into a feed-forward neural network via standard supervised learning at sufficient scale—akin to the quote famously attributed to José Raúl Capablanca, world chess champion from 1921 to 1927: “I see only one move ahead, but it is always the correct one.”
We make the following main contributions:
We distill an approximation of Stockfish 16 into a neural predictor that generalizes well to novel board states.
We construct a policy from our neural predictor and show that it plays chess at grandmaster level (Lichess blitz Elo 2895) against humans and succcessfully solves many challenging chess puzzles (up to Elo 2800). To the best of our knowledge this is currently the strongest chess engine without explicit search.
We perform ablations of the model size and data set size, showing that robust generalization and strong chess play only arise at sufficient scale.
2 Methods
@@
We now provide details on the dataset creation, the predictors and policies, and the evaluation (see Figure 1).
2.1 Data
@@
To construct a dataset for supervised training we download 10 million games from Lichess (lichess.org) from February 2023. We extract all board states from these games and estimate the state-value for each state with Stockfish 16 using a time limit of 50ms per board (unbounded depth and level). The value of a state is the win percentage estimated by Stockfish, lying between and .111 Stockfish returns a score in centipawns that we convert to the win percentage with the standard formula from https://lichess.org/page/accuracy. We also use Stockfish to estimate action-values for all legal actions in each state. Here we use a time limit of 50ms per state-action pair (unbounded depth and max skill level), which corresponds to an oracle Lichess blitz Elo of 2713 (see Section 3.1). The action-values (win percentages) also determine the oracle best action :
We rarely get time-outs when computing action-values via Stockfish, in which case we cannot determine the best action for a board state and drop the corresponding record from the behavioral cloning training set (see Table A1). Since we train on individual boards and not whole games we randomly shuffle the dataset after annotation.
@@
For our largest training dataset, based on M games, this results in B action-value estimates (or M state-value estimates and best oracle actions) to train on. To create test datasets we follow the same annotation procedure, but on k games downloaded from a different month (March 2023, M action-value estimates, k state-value estimates and best oracle actions). Since there is only a small number of early-game board-states and players often play popular openings, this i.i.d. test set contains of boards that are also in the training set. We do not remove them, as doing so would introduce distributional shift and skew test-set metrics. Finally, we also create a puzzle test set, following the procedure in Carlini (2023), consisting of k challenging board states that come with a correct sequence of moves to solve the puzzle, which we compare against in our puzzle set accuracy evaluation. Only of the puzzle set boards appear in the training set (i.e., the initial board states, not complete solution sequences). Since evaluation of puzzle solutions is slow, we use a subset of k puzzles in some of our evaluations ( overlap with training set).
@@
The predictors we train are discrete discriminators (classifiers), therefore we convert win-percentages (the ground-truth state- or action-values) into discrete “classes” via binning: we divide the interval between and uniformly into bins (non-overlapping sub-intervals) and assign a one-hot code to each bin . If not mentioned otherwise, . For our behavioral cloning experiments we train to predict oracle actions directly which are already discrete. We perform ablations for the number of bins in Section 3.4.
2.2 Model
For all our predictors we use a modern decoder-only transformer backbone (Vaswani et al., 2017; Touvron et al., 2023a, b) to parameterize a discrete probability distribution by normalizing the transformer’s outputs with a -softmax layer. The model thus outputs probabilities. The context-size is for action-value prediction, and for state-value prediction and behavioral cloning (see ‘Tokenization’ below). The output size is (the number of bins) for action- and state-value prediction and (the number of all possible legal actions) for behavioral cloning. We use learned positional encodings (Gehring et al., 2017) as the length of the input sequences is constant. Our largest model has roughly million parameters. We provide all details for the model-size ablations in Section 3.3.
@@ 2.2 Model
Board states are encoded as FEN strings which we convert to fixed-length strings of characters where the ASCII-code of each character is one token. A FEN string is a description of all pieces on the board, whose turn it is, the castling availability for both players, a potential en passant target, a half-move clock and a full-move counter. We essentially take any variable-length field in the FEN string, and convert it into a fixed-length sub-string by padding with ‘.’ if needed. We never flip the board; the FEN string always starts at rank 1, even when it is the black’s turn. We store the actions in UCI notation (e.g., ‘e2e4’ for the well-known white opening move). To tokenize them we determine all possible legal actions across games, which is , sort them alphanumerically (case-sensitive), and take the action’s index as the token, meaning actions are always described by a single token (all details in Section A.1).
@@
Predictors are trained by minimizing cross-entropy loss (i.e., -loss) via mini-batch based stochastic gradient descent using Adam (Kingma and Ba, 2015). We train for million steps, which corresponds to epochs for a batch size of with B data points (cf. Table A1). The target labels are either bin-indices in the case of state- or action-value prediction (see Section 2.1) or action indices for behavioral cloning; using a one-hot encoding in all cases (details in Sections A.2 and A.3).
@@ 2.3 Predictors and Policies
Our predictors are discrete distributions parameterized by neural networks that take a tokenized input and output a predictive distribution over discrete labels . Depending on the prediction-target we distinguish between three tasks (see Figure 1 for an overview).
@@
The target label is the bin into which the ground-truth action-value estimate falls. The input to the predictor is the concatenation of tokenized state and action. The loss for a single data point is:
where is the number of bins and is a function that computes the (one-hot) bin-index of value . To use the predictor in a policy, we evaluate the predictor for all legal actions in the current state and pick the action with maximal expected action-value:
@@
This straightforwardly gives a policy that picks the highest-probability action:
@@ 2.4 Evaluation
We use the following evaluation metrics to compare our models against each other and/or measure training progress. The first two metrics evaluate the predictors only; the second two evaluate the policies constructed from our predictors.
Action-accuracy
The test set percentage where the predictor policy picks the ground-truth best action: .
Action-ranking (Kendall’s )
The average Kendall rank correlation (a standard statistical test) across the test set, quantifying the correlation of the predicted actions with the ground-truth ranking by Stockfish in each state, ranging from -1 (exact inverse order) to 1 (exact same order) and 0 being no correlation. The predictor ranking is given by , , and , respectively, for all legal actions. The ground-truth ranking is given by Stockfish’s action-values for all legal actions.
Puzzle-accuracy
We evaluate our policies on their capability of solving puzzles from a collection of Lichess puzzles that are rated by Elo difficulty from to , calculated by Lichess based on how often each puzzle has been solved correctly. We use puzzle-accuracy as the percentage of puzzles where the policy’s action-sequence exactly matches the known solution action-sequence. For our main puzzle result in Section 3.2 we use k puzzles to report puzzle-accuracy, otherwise we use the first k puzzles to speed up evaluation.
Lichess Elo | Accuracy (%) | |||||||
Agent | Search | Input | Tournament Elo | vs. Bots | vs. Humans | Puzzles | Actions | Kendall’s |
9M Transformer (ours) | FEN | 2007 () | 2054 | - | 85.5 | 64.2 | 0.269 | |
136M Transformer (ours) | FEN | 2224 () | 2156 | - | 92.1 | 68.5 | 0.295 | |
270M Transformer (ours) | FEN | 2299 () | 2299 | 2895 | 93.5 | 69.4 | 0.300 | |
GPT-3.5-turbo-instruct | PGN | - | 1755 | - | 66.5 | - | - | |
AlphaZero (policy net only) | PGN | 1620 () | - | - | 61.0 | - | - | |
AlphaZero (value net only) | PGN | 1853 () | - | - | 82.1 | - | - | |
AlphaZero (400 MCTS simulations) | ✓ | PGN | 2502 () | - | - | 95.8 | - | - |
Stockfish 16 (0.05s) [oracle] | ✓ | FEN | 2706 () | 2713 | - | 99.1 | 100.0 | 1.000 |
@@
We evaluate the playing strength (measured as an Elo rating) of the predictor policies in two different ways: (i) we play Blitz games on Lichess against either only humans or only bots, and (ii) we run an internal tournament between all the agents from Table 1 except for GPT-3.5-turbo-instruct. We play 400 games per pair of agent, yielding 8400 games in total, and compute Elo ratings with BayesElo (Coulom, 2008), with the default confidence parameter of . We anchor the relative BayesElo values to the Lichess ELO vs. bots of our 270M model. For the game-playing strength evaluations only (i.e., not for determining the puzzle accuracy) we use a softmax policy for the first 5 full-moves, instead of the policy described earlier, with a low temperature of for the value or action-value functions, for the action functions (like the policy network of AlphaZero), and for the visit counts used in the full version of AlphaZero. This renders the policies stochastic to both create variety in games, and prevents simple exploits via repeated play.
2.5 Baselines
We compare the performance of our models against Stockfish 16 (with a time limit of 0.05s per legal move, i.e., the oracle used to generate our dataset), three variants of AlphaZero (Silver et al., 2017): (i) the original with 400 MCTS simulations, (ii) only the policy network, and (iii) only value network (where (ii) and (iii) perform no additional search), and the GPT-3.5-turbo-instruct from Carlini (2023). AlphaZero’s networks have M parameters and are trained on M games (details in Schrittwieser et al. (2020)). Note that these baselines have access to the whole game history (via the PGN), in contrast to our models that only observe the current game state (which contains very limited historical information via the FEN). This helps the baseline policies for instance to easily deal with threefold repetition (games are drawn if the same board state is appears three times throughout the game), which requires a workaround for us (described in Section 5). Moreover, GPT-3.5-turbo-instruct also requires whole games encoded via PGN to reduce hallucinations according to Carlini (2023), who also finds that GPT-4 struggles to play full games without making illegal moves, so we do not compare against GPT-4.
3 Results

### v2

# Amortized Planning with Large-Scale Transformers: A Case Study on Chess

@@ Abstract
This paper uses chess, a landmark planning problem in AI, to assess transformers’ performance on a planning task where memorization is futile — even at a large scale. To this end, we release ChessBench, a large-scale benchmark dataset of million chess games with legal move and value annotations ( billion data points) provided by Stockfish 16, the state-of-the-art chess engine. We train transformers with up to million parameters on ChessBench via supervised learning and perform extensive ablations to assess the impact of dataset size, model size, architecture type, and different prediction targets (state-values, action-values, and behavioral cloning). Our largest models learn to predict action-values for novel boards quite accurately, implying highly non-trivial generalization. Despite performing no explicit search, our resulting chess policy solves challenging chess puzzles and achieves a surprisingly strong Lichess blitz Elo of 2895 against humans (grandmaster level). We also compare to Leela Chess Zero and AlphaZero (trained without supervision via self-play) with and without search. We show that, although a remarkably good approximation of Stockfish’s search-based algorithm can be distilled into large-scale transformers via supervised learning, perfect distillation is still beyond reach, thus making ChessBench well-suited for future research.
@@ 1 Introduction
The ability to plan ahead and reason about long-term consequences of actions is a hallmark of rationality and human intelligence. Its replication in artificial systems has been a central goal of AI research since the field’s inception. Perhaps the historically most famous planning problem in AI is chess, where naive search is computationally intractable and brute-force memorization is futile — even at scale. Conceptually, the ability to plan and reason about consequences of actions is implemented via search algorithms and value computation [1, 2]. To implement such algorithms at scale, feed-forward neural networks have been playing an increasingly important role. For instance, DQN [3] and AlphaGo [4] popularized the use of neural value estimators to scale RL and (Monte Carlo) tree search. Similarly, today’s strongest chess engines also employ a combination of search and amortized neural value estimation. For example, the state-of-the-art Leela Chess Zero [5, 6], which builds upon AlphaZero [7], augments Monte Carlo tree search with neural value predictions. Likewise, Stockfish 16, currently the strongest chess engine, uses efficiently updatable neural network evaluation [8], a highly specialized neural architecture to obtain fast evaluations. It is trained on value estimates of an earlier Stockfish version that uses human chess heuristics. All these chess engines tweak value-estimation networks towards fast evaluations that are to be combined with search over many possible future move sequences. As a result, the value estimates themselves may not necessarily be optimal, which leaves open the question of how far amortized planning with modern neural networks can be pushed with scale and whether the resulting “searchless” engines can match engines that use search at test time, at least in principle (putting aside computational efficiency).
To scientifically address this question, we create ChessBench (https://github.com/google-deepmind/searchless_chess), a large-scale chess dataset created from 10 million human games that we annotate with Stockfish 16. We use ChessBench to train transformers of up to million parameters via supervised learning to predict action-values given a board-state. We also construct searchless chess policies from these predictors, where the playing strength depends entirely on the quality of the value predictions. We find that predictions by our largest trained models generalize well and non-trivially to novel board states. The resulting policies are capable of solving challenging chess puzzles and playing chess at a high level (grandmaster) against humans. Due to the combinatorial explosion of chess board states, virtually every new game involves many board states that were never seen during training (see Figure 1). This rules out memorization as a possible explanation and suggests that an approximate version of Stockfish’s (search-based) value estimation algorithm can indeed be distilled into large transformers. Nonetheless, we also find that the performance gap cannot be fully closed, which may indicate that architectural innovations (or better optimization, data augmentation, etc.) might be needed instead of even larger scale.
Our main contributions are:
We introduce ChessBench, a large-scale benchmark dataset for chess, consisting of M board states (from M games on lichess.org) annotated via Stockfish 16 with state-values and best-action, as well as B action-values for all legal actions in each board state (corresponding to roughly 8864 days of unparallelized Stockfish evaluation time).
We train transformers of up to M parameters to predict action-values via supervised learning. The largest models generalize well to novel boards, enable play at grandmaster level (Lichess blitz Elo of 2895) against humans, and solve chess puzzles with dificulty ratings up to a Lichess Puzzle Elo of – without using explicit search at test time.
We perform extensive ablations, including model- and dataset size, network architecture, action-value vs. state-value vs. behavioral cloning, and various hyper-parameters.
We open source our ChessBench dataset, our model weights, and all training and evaluation code at https://github.com/google-deepmind/searchless_chess, and provide a series of benchmark results through our models, ablations, and comparisons with Stockfish 16, Leela Chess Zero, and AlphaZero, and their searchless policy/value networks.
2 Methodology
@@
We now describe the dataset creation, the neural predictors and how to construct policies from them, and our evaluation methodology (see Figure 1 for an overview; full details in Appendix A).
2.1 ChessBench Dataset
@@
To construct a training dataset for supervised learning we downloaded 10 million games from Lichess (lichess.org) from February 2023111Lichess games and puzzles are released under the Creative Commons CC0 license.. We extract all board states from these games and estimate the state-value for each state with Stockfish 16 (our “value oracle”) using a time limit of 50ms per board (unbounded depth and maximum skill level). The value of a state is the win percentage estimated by Stockfish, lying between and .222 Stockfish returns a score in centipawns that we convert to the win percentage with the standard formula from https://lichess.org/page/accuracy. Note that Stockfish does not provide a centipawn score when it detects a mate-in-, so we map all of these cases to a win percentage of %. We also use Stockfish to estimate action-values for all legal actions in each state. Here we use a time limit of 50ms per state-action pair (unbounded depth and max skill level), which corresponds to a Lichess blitz Elo of 2713 for our action-value oracle (see Section 3.1). The action-values (win percentages) also determine the oracle best action :
In the case where multiple moves are tied in value, we pick a maximizing action arbitrarily. Since we train on individual boards and not whole games we shuffle the dataset after annotation. For our largest training dataset, based on M games, we thus obtain B action-value estimates (or M state-value estimates and oracle best-actions; cf. Table A1) to train on (corresponding to roughly days of unparallelized Stockfish 16 evaluation time given the limit of 50ms per move).
@@
To create our test dataset we follow the same annotation procedure, but on K games downloaded from a different month (March 2023), resulting in M action-value estimates (or K state-value estimates and oracle best actions). Since there is only a small number of early-game board states and players often play popular openings, of boards in this i.i.d. test are also in the training set. We deliberately chose not remove them, to not introduce a distributional shift and skew test set metrics. We also create a puzzle test set, following Carlini [9], consisting of K challenging board states that have an Elo rating and a sequence of moves to solve the puzzle. Only of the puzzle boards appear in the training set (i.e., the initial board states, not the complete solution sequences).
2.2 Data Preprocessing and Training
@@
The predictors we train are discrete discriminators (i.e., classifiers). Therefore we convert the win percentages (i.e., the ground truth state- or action-values) into discrete “classes” via binning, akin to distributional RL [10]. We divide the interval from to uniformly into bins (non-overlapping sub-intervals) and assign a one-hot code to each bin . If not mentioned otherwise, . For our behavioral cloning experiments we train to directly predict the oracle actions, which are already discrete. We ablate the number of bins in Section B.1, and we investigate losses that do (i.e., the loss) and do not (i.e., the HL-Gauss and L2 loss) maintain semantic interconnectedness of the labels (i.e., the dependence between the classes) in Table 2.
@@
A FEN [11] string is a standard string-based description of a chess position. It consists of a board state, the current side to move, the castling availability for both players, a potential en passant target, a half-move clock and a full-move counter, all represented in a single ASCII string. Our tokenization uses this representation, except that we flatten any uses of run-length encoding to obtain a fixed length (77) tokenized representation. Actions are stored in UCI notation [12], for example ‘e2e4’ for the popular opening move for white. To tokenize we use the index of a move into a lexiographically sorted array of the 1968 possible UCI encoded actions (see Section A.1 for more details). Note that this representation technically makes the game non-Markovian, because FENs do not contain the move history and therefore, cannot capture all information for rules such as drawing by threefold repetition (drawing because the same board occurs three times).
Network inputs and outputs
For all our predictors we use a modern decoder-only transformer backbone [13, 14, 15] to parameterize a categorical distribution by normalizing the transformer’s outputs with a layer. The models thus output probabilities. The context size is for action-value prediction and for state-value prediction and behavioral cloning (see ‘Tokenization’ above). The output size is (i.e., the number of bins) for action- and state-value prediction and (the number of all possible legal actions) for behavioral cloning. We use learned positional encodings [16] as the length of the input sequences is constant. Our largest model has million parameters, i.e., layers, heads, and an embedding dimension of (full details in Section A.2).
@@
We train our predictors by minimizing the cross-entropy loss via mini-batch stochastic gradient descent using Adam [17]. The target labels are either bin-indices for state- or action-value prediction or action indices for behavioral cloning. For state- and action-value prediction, we additionally apply label smoothing via the HL-Gauss loss [18], using a Gaussian smoothing distribution with the mean given by the label and a standard deviation of as recommended by Farebrother et al. [19]. Thus, the labels are no longer one-hot but multi-hot encoded (see Figure 3 in [19] for an overview). We ablate the loss function, i.e., HL-Gauss vs. cross-entropy vs. MSE, in Section 3.4. We train for million steps, which corresponds to epochs for a batch size of with B data points (cf. Table A1; details in Sections A.2 and A.3).
@@ 2.3 Predictors and Policies
Our predictors are categorical distributions parameterized by neural networks that take a tokenized input and output a predictive distribution over discrete labels . Depending on the prediction target we distinguish between three tasks (see Figure 1 for an overview).
@@
The target label is the bin into which the ground-truth action-value estimate falls. The input to the predictor is the concatenation of tokenized state and action. The loss for a single input data point is:
where is the number of bins and is the function that computes (the smoothed version of label ), i.e., a smooth categorical distribution using HL-Gauss [18] rather than the standard one-hot procedure. To use the predictor in a policy, we evaluate the predictor for all legal actions in the current state and pick the action with maximal expected action-value:
@@
This gives a policy that picks the highest-probability action:
@@ 2.4 Evaluation
We use the following metrics to evaluate our models and/or measure training progress.
Action accuracy
The test set percentage where the policy picks the best action: .
Action ranking (Kendall’s )
The average test set rank correlation [20] of the predicted action distribution with the ground truth given by Stockfish. Kendall’s ranges from -1 (exact inverse order) to 1 (exact same order), with 0 indicating no rank correlation. The ranking is obtained by evaluating for all legal actions , , or , in the case of AV, SV, or BC prediction, respectively. The ground-truth ranking is given by evaluating for all legal actions.
Puzzle accuracy
We evaluate our policies on their capability of solving puzzles from a collection of Lichess puzzles that are rated by Elo difficulty from to based on how often each puzzle has been solved correctly. We use puzzle accuracy as the percentage of puzzles where the policy’s action sequence exactly matches the entire solution action sequence. For our main results in Sections 3.1 and 3.2 we use K puzzles; otherwise, we use the first K puzzles to speed up evaluation.
@@
We evaluate the playing strength (measured as an Elo rating) of the policies in two ways: (i) we run an internal tournament between all the agents from Table 1 except for GPT-3.5-turbo-instruct, and (ii) we play Blitz games on Lichess against either only humans or only bots. For the tournament we play games per agent pair, yielding K games in total, and compute the Elo with BayesElo [21] using the default confidence parameter of . To ensure variability in the games, we use the openings from the Encyclopaedia of Chess Openings (ECO) [22]. When playing on Lichess, we use a policy with a low temperature of for the first five full-moves instead of the policy to create variety in games and prevent simple exploits via repeated play. We anchor the relative BayesElo values to the Lichess Elo vs. bots of our largest (M) model.
2.5 Engine Comparisons
We compare the performance of our models against the following engines:
Stockfish 16
We consider two variants: (i) a ms time limit per legal move (i.e., the dataset oracle), and (ii) a s limit per board, which is roughly the amount of time the first variant (i) takes on average per board (i.e., there are roughly legal moves per board on average).
AlphaZero
Leela Chess Zero
We consider three variants: (i) with 400 MCTS simulations, (ii) only the policy network, and (iii) only the value network (where (ii) and (iii) perform no additional search). We use the T82 network, which is a convolutional network with 768 filters, 15 blocks, and mish activations [24] – the largest network available on the official Leela Chess Zero website. At the time of writing, the precise network architecture, training data, and training protocol for the T82 network were not reported on the Leela Chess Zero website or other source. Concurrently to our paper, Monroe and Leela Chess Zero Team [6] published a rigorous tech report which now provides many of these details about for a slightly different (state-of-the-art) Leela Zero architecture called ChessFormer, and includes a comparison against our vanilla networks.
GPT-3.5-turbo-instruct
We follow Carlini [9] and encode the entire game with the Portable Game Notation (PGN) [11] to reduce hallucinations. Since Carlini [9] found that (at that time) GPT-4 struggled to play full games without making illegal moves, we do not consider GPT-4.
In contrast to our models, all of the above engines rely on the entire game’s history (via the PGN; we use the FEN, which only contains the game’s current state and very limited historical information). Observing the full game’s history helps, for instance, detecting and preventing draws from threefold repetition (games are drawn if the same board state appears three times throughout the game). Our engines require a workaround to deal with this problem (described in Section 4).
A direct comparison between all engines comes with a lot of caveats since some engines use the game history, some have very different training protocols (i.e., RL via self-play instead of supervised learning), and some use search at test time. We show these comparisons to situate the performance of our models within the wider landscape, but emphasize that some conclusions can only be drawn within our family of models and the corresponding ablations that keep all other factors fixed.
3 Case-Study Results